The magnitude of the gravitational force between two point masses separated by a distance of 66.3 cm is 11.2 μN. If the masses are moved so that they are separated by a distance of 42.9 cm, what is the magnitude of the gravitational force they exert on each other after they have been moved?

  • 11.2 N
  • 4.69 N
  • 7.25 N
  • 17.3 N
  • 26.8 N
Answer #1

here,

the magnitude of gravitational force between masses , Fg = G * m1 * m2 /r^2

when r1 = 66.3 cm , F1 = 11.2 uN

when r2 = 42.9 cm , let the force be F2

we can write

F2 /F1 = ( G * m1 * m2 /r2^2) /(G * m1 * m2 /r1^2)

F2 /F1 = ( r1 /r2)^2

F2 /11.2 = (66.3/42.9)^2 uN

F2 = 26.8 uN

the magnitude of the gravitational force they exert on each other after they have been moved is 26.8 uN